Jim's weekly pay is two thirds of Alicias's. Together they earn 600 dollars per week. What is each person's weekly pay?
Answer by jainenderkapoor(61)

Let us suppose that Alicia's Weekly Pay is $ x
Then Jim's weekly pay becomes = 2/3 x
So now we can write a simple linear equation
x + 2/3 x = 600
x(1 + 2/3) = 600
x 5/3 = 600
x = 600 * 3/5
x = 1800/5 = 360
Therefore Alicia's Weekly pay is $ 360
and Jim's Weekly pay is 2/3 * 360 = $ 240
Hope the solution is clear to you.
In case you need any other help on any topic, you are welcome to contaft me.
I am available for ONLINE TUTORING at very economical rates.
My mail id is kapoorjai1@rediffmail.com.